OpenStudy (erniellyd28):

a student scores 71 out of 100 points on a test. if the maximum score on the next test is also 100 what score does the student need to maintain an average of 78?

OpenStudy (littleelf):

What do u think

OpenStudy (littleelf):

\[e\]

OpenStudy (littleelf):

opps

OpenStudy (littleelf):

Ok there u go

OpenStudy (littleelf):

actully

OpenStudy (littleelf):

wait

OpenStudy (littleelf):

you have: (74 + x)/2 = 80 multiply both sides of this equation by 2 to get: 74 + x = 160 subtract 74 from both sides of this equation to get: x = 160 - 74 which becomes: x = 86 He has to get 86 on the second test to average 80 on both tests. Assuming he gets that, then his grade average would be calculated as: (74 + 86) / 2 = 160/2 = 80.

OpenStudy (erniellyd28):

i was gonna say 80 but i wasnt 100% sure

OpenStudy (littleelf):

Lol

OpenStudy (littleelf):

Make sure you medal whoever helps you :)

OpenStudy (erniellyd28):

its like you knew what i was thinking and or gonna say lol

OpenStudy (littleelf):

xD

OpenStudy (erniellyd28):

i always do :-)

OpenStudy (littleelf):

ty

OpenStudy (erniellyd28):

i checked to make sure it was 80 and it was wrong the computer said it was 85

OpenStudy (littleelf):

Ummm

OpenStudy (littleelf):

Oh sorry :(

OpenStudy (erniellyd28):

its ok its all good :-)

OpenStudy (erniellyd28):

dont trip potato chip lol

OpenStudy (littleelf):

lol xD

OpenStudy (erniellyd28):

but why does the comp say its 85